Question 1 - Selection of an activity base Fantastic Sweepers of New England manufactures replacement rotary sweeper brooms fThe number of workdays varies from month to month due to the number of weekdays and holidays, as well as vacation days and siQuestion 2 - Mixed cost analysis and prediction The number of X-rays taken and X-ray costs over the first nine months of 2019Question 3 - Mixed cost analysis and relevant ranges The Roakes Company is interested in developing a cost equation to estimaDecember 18.000 $75.000 The company leases all of its manufacturing equipment. The lease arrangement calls for a flat monthlyplease solve neat and clean handwriting.
Answer #1
Q1 Arrange Units Produced in Ascending Order
Units Produced Work days Labor Cost
69 23 $4,320
71 14 $2,688
76 20 $3,456
77 16 $2,784
102 23 $4,128
108 22 $4,032
109 19 $3,648
115 21 $3,840
127 21 $3,840
From the Scatter diagram , it is clear that Labor Cost is correlated with Work days.
Janitorial Workdays should be used as activity base
Q2 X Y
X Ray Taken X Ray Cost
3000 $17,000
3750 $18,000
4250 $20,000
4500 $22,000
5000 $23,000
5500 $24,000
5750 $26,000
6250 $28,000
7000 $27,500
High Low Method
X Y
X ray taken X ray Cost
High 7000 $27,500
Low 3000 $17,000
Y=a+bX
27500=a+b*7000 ….Equation(1)
17000=a+b*3000 ….Equation(2)
Subtraction Equation(2) from Equation(1)
b*4000= $10,500
b=10500/4000                  2.6250
17000=a+b*3000
17000=a+2.6250*3000
a=17000-(2.6250*3000)= $9,125
COST EQUATION:
X-Ray Cost =$9125+2.63*(X ray taken)
Y=$9125+2.63*X
(b) X ray taken X= 4600
Y=$9125+2.6250*X= $21,200.00
X ray Cost = $21,200.00
Click "Data"
Click "Data Analysis"
Select "Regression" tool
Click"OK"
Input y_Range (X-Ray Costs)
Input x_Range (X ray taken)
Click "OK"
Refer To Summary Output
Intercept 7539.22
X Variable 3.06
Cost Equation:
Y=7539.22+3.06X

UNITS PRODUCED VS LABOR COST;X-axis-Units Y-Axis-Cost $5,000 $4,500 $4,000 $3,500 $3,000 $2,500 $2,000 $1,500 $1,000 $500 20WORKDAYS VS LABOR COST:X axis: Workdays, Y axis-Labor Cost $5,000 $4,500 $4,000 $3,500 $3,000 $2,500 $2,000 $1,500 $1,000 101 SUMMARY OUTPUT 3 Regression Statistics 4 Multiple F 0.973026 5 R Square 0.946779 6 Adjustedi 0.939176 7 Standard I 978.7662Number of X ray taken Vs Actual Cost and Cost as per Equation $35,000 $30,000 $25,000 $20,000 9515,000 $10,000 $5,000 1000 20Y1 Cost as X Ray per X Ray Taken Cost(Actual) Equation Y=7539.22+3.06X 3000 $17,000 $16,719 3750 $18,000 $19,014 4250 $20,000